Determine the values of sin(2a) and tan(2a), given tan(a) = -7/24 and Pi over 2 ≤ 0 ≤ Pi
1
Expert's answer
2020-10-25T18:55:24-0400

"\displaystyle
\tan{a} = \frac{-7}{24}\\

\tan(\pi - a) = \frac{7}{24}\\

a = \pi - \arctan\left(\frac{7}{24}\right)\\

\begin{aligned}
\sin(2a) &= 2\cos{a}\sin{a} 
\\&= 2\cos\left(\pi - \arctan\left(\frac{7}{24}\right)\right)\sin\left(\pi - \arctan\left(\frac{7}{24}\right)\right) 
\\&= -2\cos\left(\arctan\left(\frac{7}{24}\right)\right) \sin\left(\arctan\left(\frac{7}{24}\right)\right) 
\end{aligned}\\

\textsf{Let}\,y = \cos(\arctan(x))\\

y = \cos{u}, x = \tan{u}\\

1 + x^2 = \sec^2{u}\\

\cos{u} = \frac{1}{\sqrt{1 + x^2}}\\

y = \cos{u} = \frac{1}{\sqrt{1 + x^2}}\\

\begin{aligned}
\sin{u} &= \sqrt{1 - \cos^2{u}} 
\\&= \sqrt{1 - \left(\frac{1}{\sqrt{1 + x^2}}\right)^2} = \frac{x}{\sqrt{1 + x^2}} = \sin(\arctan(x))
\end{aligned}\\

\begin{aligned}
\sin(2a) &= -2 \times \frac{1}{\sqrt{1 + \frac{7^2}{24^2}}} \times \frac{7/24}{\sqrt{1 + \frac{7^2}{24^2}}} 
\\&= \frac{-14 \times 24}{(\sqrt{7^2 + 24^2})(\sqrt{7^2 + 24^2})}
\\&= \frac{-14 \times 24}{7^2 + 24^2} = -\frac{336}{625}
\end{aligned}\\

\begin{aligned}
\tan(2a) &= \frac{2\tan{a}}{1 - \tan^2{a}} 
\\&= \frac{2 \times -7/24}{1 - \frac{7^2}{24^2}} = -\frac{7}{12} \times \frac{24^2}{24^2 - 7^2}
\\&= \frac{-14 \times 24}{24^2 - 7^2} = \frac{-336}{527}
\end{aligned}\\"